### v1

# Supervised Learning Models for Early Detection of Albuminuria Risk in Type-2 Diabetes Mellitus Patients

Abstract
Diabetes, especially T2DM, continues to be a significant health problem. One of the major concerns associated with diabetes is the development of its complications. Diabetic nephropathy, one of the chronic complication of diabetes, adversely affects the kidneys, leading to kidney damage. Diagnosing diabetic nephropathy involves considering various criteria, one of which is the presence of a pathologically significant quantity of albumin in urine, known as albuminuria. Thus, early prediction of albuminuria in diabetic patients holds the potential for timely preventive measures. This study aimed to develop a supervised learning model to predict the risk of developing albuminuria in T2DM patients. The selected supervised learning algorithms included Naïve Bayes, Support Vector Machine (SVM), decision tree, random forest, AdaBoost, XGBoost, and Multi-Layer Perceptron (MLP). Our private dataset, comprising 184 entries of diabetes complications risk factors, was used to train the algorithms. It consisted of 10 attributes as features and 1 attribute as the target (albuminuria). Upon conducting the experiments, the MLP demonstrated superior performance compared to the other algorithms. It achieved accuracy and f1-score values as high as 0.74 and 0.75, respectively, making it suitable for screening purposes in predicting albuminuria in T2DM. Nonetheless, further studies are warranted to enhance the model’s performance.
Index Terms:
diabetes, albuminuria, supervised learning, machine learning, deep learning
I Introduction
Diabetes continues to be one of the most challenging noncommunicable diseases worldwide. It is a chronic metabolic disorder characterized by high blood sugar levels caused by problems in insulin production, sensitivity of cells’ response to insulin, or both [1]. There are four types of diabetes, namely type-1, type-2, gestational type, and other types. However, type-2 diabetes (T2DM) dominates all other diabetes types [2], accounting for more than 90% of all diabetes cases. The high prevalence of T2DM is strongly associated with the unhealthy modern lifestyle, including unhealthy eating habits, smoking, obesity, and a lack of physical activity, as well as internal predisposition factors such as race and family history [3].
The predominant challenge associated with diabetes stems from the array of complications that can arise when diabetes is not adequately controlled. Among these unwanted complications, one particularly notable issue is kidney complication, which falls under the category of microvascular complications, affecting the smaller blood vessels [4, 5]. This specific complication is commonly referred to as diabetic nephropathy and accounts for approximately 14.0% of diabetes-related complications [4].
Diabetic nephropathy is considered as a type of Chronic Kidney Disease (CKD). According to the Kidney Disease Improving Global Outcomes (KDIGO) 2012 guidelines, CKD is established when there are markers of kidney damage and/or a Glomerular Filtration Rate (GFR) $<$ 60 mL/min/1.73m2 that lasts for at least $\geq$ 3 months. The kidney damage markers for CKD include the presence of pathologically high quantities of urinary albumin excretion (albuminuria), the presence of urine sediment abnormalities, structural abnormalities detected by imaging, and a history of kidney transplantation [6].
As mentioned in the preceding paragraph, the presence of albuminuria can be indicative of a kidney problem. Albumin in urine can signal an issue with the kidney filtration function. Albuminuria can be divided into two categories: microalbuminuria and macroalbuminuria. Microalbuminuria is diagnosed when the albumin-creatinine ratio is $>$ 30 mg/24h and $<$ 300 mg/24h, while macroalbuminuria is diagnosed when the albumin excretion is $>$ 300 mg/24h in a 24-hour urine collection sample [7]. As albuminuria can serve as a signal of kidney problems, it becomes essential for diabetes patients to be aware of their risk of developing this condition.
Therefore, the primary objective of this study is to develop a supervised learning model capable of predicting the risk of albuminuria development in diabetes patients, particularly those with T2DM. The primary contributions of this paper can be summarized as follows:
  • Development of a supervised model capable of predicting early albuminuria in patients with type 2 diabetes mellitus (T2DM).
  • Identification of the optimal supervised algorithm for early albuminuria detection in T2DM patients.
II Methodology
II-A Dataset
Figure 1: Dataset distribution
TABLE I: Summary Statistics
Attribute Unit Mean±$\pm$±Std Min Max
durasi_dm Year 5.168$\pm$5.012 0.000 38.000
bmi kg/m2 27.973$\pm$4.517 18.350 44.870
hdl mg/dL 48.065$\pm$10.913 22.000 88.000
ldl mg/dL 135.978$\pm$38.522 40.000 307.000
tg mg/dL 185.571$\pm$118.305 47.000 878.000
kol_tot mg/dL 207.554$\pm$42.534 105.000 347.000
gdp mg/dL 158.076$\pm$68.591 74.000 433.000
TDS mmHg 141.609$\pm$18.682 96.000 201.000
a1c % 8.204$\pm$2.134 5.300 16.300
cr mg/dL 0.782$\pm$0.409 0.300 4.810
Attribute Category Count Perc.
kid_group Normal (0) 92 0.50
Albuminuria (1) 92 0.50
In this study, we used our private dataset consisting of data on the risk of diabetes complications, which was collected from a primary healthcare facility in DKI Jakarta, Indonesia. The dataset comprises 184 records, each consisted of 10 features and 1 target variable (Table I) (Figure 1). All records are sourced from patients with T2DM. The features are all numerical, whereas the target variable is categorical. Prior to analysis, all data were carefully examined and cleaned to remove any missing values or measurement errors. To ensure the security and privacy of the medical data, all information was anonymized.
  • durasi_dm: This attribute refers to the length of time since the patient’s initial diabetes diagnosis. The duration is measured in years.
  • bmi: This attribute refers to the patient’s current Body Mass Index (BMI), which is measured in kg/m2.
  • hdl: This attribute refers to the current level of High-Density Lipoprotein (HDL) in the bloodstream, measured in mg/dL using standard laboratory methods.
  • ldl: This attribute refers to the current level of Low-Density Lipoprotein (LDL) in the bloodstream, measured in mg/dL using standard laboratory methods.
  • tg: This attribute refers to the current level of triglyceride in the bloodstream, measured in mg/dL using standard laboratory methods.
  • kol_tot: This attribute refers to the current level of total cholesterol in the bloodstream, measured in mg/dL using standard laboratory methods.
  • gdp: This attribute refers to the current level of fasting plasma glucose in the bloodstream, measured in mg/dL using standard laboratory methods.
  • TDS: This attribute refers to the current systolic blood pressure measured in mmHg using an ambulatory blood pressure device.
  • a1c: This attribute refers to the current level of HbA1c measured using standard laboratory methods.
  • cr: This attribute refers to the current level of creatinine, measured in mg/dL using standard laboratory methods.
  • kid_group: This attribute serves as the target label and describes the grouping of kidney disease. It is a categorical attribute comprising of two categories: normal and albuminuria. The determination of albuminuria label was based on the KDIGO 2012 criteria. However, instead of treating microalbuminuria and macroalbuminuria as separate categories, we classified them both under the umbrella term of ”albuminuria”.
The use of the aforementioned features was rationalized based on the complex nature of their interaction with kidney damage, as shown in Figure 2 [8, 9, 10, 11].
Figure 2: Simplified diabetic nephropathy mechanism
II-B Design of Experiment
This study aimed to evaluate the performance of supervised learning algorithms in predicting the risk of developing albuminuria in patients with T2DM patients. We evaluated several supervised learning algorithms, including 6 machine learning algorithms and 1 deep learning algorithm. The machine learning algorithms used were Naïve Bayes, Support Vector Machine (SVM), decision tree, random forest, AdaBoost, and XGBoost. Among these machine learning algorithms, random forest, AdaBoost, and XGBoost are ensemble algorithms. The deep learning algorithm employed in this study is the Multi-Layer Perceptron (MLP). Table II presents the complete experimental design used in this study. We used the scikit-learn library [12], version 1.0.2, as our primary machine learning and deep learning toolkit. Additionally, we employed the xgboost library [13], version 1.6.2, specifically designed for implementing the XGBoost algorithm.
TABLE II: Design of Experiment
Algorithm Observable Factors
Naïve Bayes Class sklearn.naive_bayes.GaussianNB
Params priors=None, var_smoothing=1e-09
SVM Class sklearn.svm.SVC
Params C=1.0, kernel=’rbf’, max_iter=-1
Decision Tree Class sklearn.tree.DecisionTreeClassifier
Params criterion=’gini’, max_depth=None
Random Forest Class sklearn.ensemble.RandomForestClassifier
Params n_estimators=100, criterion=’gini’,
max_depth=None
AdaBoost Class sklearn.ensemble.AdaBoostClassifier
Params n_estimators=50, learning_rate=1.0,
algorithm=’SAMME.R’
XGBoost Class xgboost.XGBClassifier
Params n_estimators=2, max_depth=1,
learning_rate=1
MLP Class sklearn.neural_network.MLPClassifier
Params hidden_layer_sizes=(100,),
learning_rate_init=3e-3,
max_iter=200
The dataset is split into training and test datasets using the train_test_split function provided by scikit-learn. Since the dataset size is relatively small, we opted for a train-test ratio of 0.75:0.25.
II-C Evaluation Strategy
We used precision (1), recall (2), accuracy (3), and f1-score (4) as the evaluation metrics for our study. A competent model is expected to exhibit high values for precision, recall, accuracy, and f1-score.
$precision=\frac{TP}{TP+FP}$ (1)
$recall=\frac{TP}{TP+FN}$ (2)
$accuracy=\frac{TP+TN}{TP+FP+FN+TN}$ (3)
$F1\textit{-}score=\frac{2\times precision\times recall}{precision+recall}$ (4)
II-D Ethics Approval
This study has been ethically approved by the Health Ethics Committee of Cipto Mangunkusumo Hospital, Faculty of Medicine Universitas Indonesia, Jakarta, Indonesia number KET-246/UN2.F1/ETIK/PPM.00.02/2022.
III Result
Figure 3: Distribution of the train-test dataset
Figure 4: Error analysis
Figure 3 shows the distribution of the train and test datasets used in our study. The test ratio is set at 0.25, resulting in 138 records for the train dataset and 46 records for the test dataset. As depicted in Figure 3, the dataset exhibits a relatively uniform distribution. Although several data points appear to be potential outliers, we made the decision to retain them deliberately to introduce bias to the learning algorithms and thus reduce the risk of overfitting.
We train each supervised learning algorithm with its defined parameters, as shown in Table II, using the training dataset, which results in trained models. These models are then tested using the test dataset, producing predicted labels. The performance of the models is evaluated by comparing these predicted labels with their corresponding true labels, which serve as the ground truth.
The model evaluation results are shown in Table III. As depicted in the table, the machine learning algorithms did not yield satisfactory results. Among the various machine learning algorithms experimented with, Naïve Bayes outperformed the others. It demonstrated an accuracy of up to 0.65 when predicting the test dataset. In contrast, the remaining machine learning algorithms exhibited prediction accuracies of no more than 0.5. Even the ensemble models failed to surpass Naïve Bayes in predicting the test dataset. However, it’s worth noting that none of these machine learning algorithms achieved an f1-score higher than 0.5, suggesting that the model is insufficient for predicting the risk of albuminuria in T2DM patients.
TABLE III: Classification Report
Algorithm Precision Recall Accuracy F1-Score
Naïve Bayes 0.67 0.4 0.65 0.50
SVM 0.50 0.70 0.57 0.58
Decision Tree 0.45 0.50 0.52 0.48
Random Forest 0.50 0.60 0.57 0.55
AdaBoost 0.44 0.40 0.52 0.42
XGBoost 0.46 0.55 0.52 0.50
MLP 0.68 0.75 0.74 0.71
The superior results were obtained from the deep learning algorithm, specifically the Multi-Layer Perceptron (MLP), which achieved an accuracy and f1-score of 0.74 and 0.71, respectively. This algorithm outperformed the machine learning algorithms, which only achieved accuracy and f1-scores of up to 0.65 and 0.55, respectively. Additionally, the MLP algorithm exhibited the highest precision and recall scores compared to the other algorithms, scoring 0.68 and 0.75, respectively. This result indicates that the algorithm might be acceptable for predicting the risk of albuminuria among T2DM patients. However, further improvements are needed, particularly in terms of the dataset size and variety, to achieve better results.
To gain a better understanding of the model evaluation results, we conducted a visual error analysis on the prediction outcomes of the MLP model. To facilitate visualization, we used the Principal Component Analysis (PCA) method to reduce the features from 10 to 2 dimensions. Subsequently, we used square and triangle markers to represent the normal and albuminuria labels, respectively, while using red and green colors to indicate false and true predictions, respectively. The visualization of the model evaluation results can be observed in Figure 4.
As shown in Figure 4, the false predictions could have either a normal or albuminuria label. However, the interesting point revealed by the visualization in the figure is that the falsely predicted labels are spread out but relatively close to the adjacent cluster that forms the true predictions. This indicates that the data characteristics of ’normal’ and ’albuminuria’ at some points have little difference, which might cause the algorithm to experience difficulty in creating separate boundaries between the labels, leading to false predictions and resulting in lower accuracy. This phenomenon may be explained by the nature of the patient data.
For several patients, the risk of developing albuminuria might not be strongly correlated with the features in the dataset. For example, there could be a patient with uncontrolled diabetes, indicated by high blood glucose and high lipid profile, but not developing albuminuria, while another patient with normal glucose levels and normal lipid profile developing albuminuria. This complexity arises because the human body is complex, and the risk of developing a disease may be influenced by multiple risk factors that are not apparent in the dataset. Therefore, one possible solution to improve the model’s accuracy is to increase the dataset size and variety, allowing the learning algorithms to better understand the complex patterns present in such data.
Despite the complex nature of the patient data, the outperforming performance of the MLP algorithm might be beneficial due to its architecture. The MLP consists of several to even thousands of hidden layers. The MLP uses the backpropagation algorithm to update its weights based on the learned data [14]. This enables the MLP to solve complex problems relatively easily compared to other traditional machine learning algorithms when dealing with such complex data.
IV Conclusion
We have developed a supervised learning model to predict the risk of developing albuminuria in patients with T2DM. Among the various supervised learning models examined in this study, the MLP algorithm demonstrated superior performance in terms of precision, recall, accuracy, and f1-score. Specifically, the algorithm achieved values of 0.68, 0.75, 0.74, and 0.71 for precision, recall, accuracy, and f1-score, respectively. To further enhance the model’s performance, we recommend augmenting the dataset with additional data to increase its size and diversity. Additionally, we propose conducting further research into the utilization of deep learning algorithms like MLP to effectively handle the complexities inherent in patient data.